OpenStudy (anonymous):

Solve: −2(x − 1) = −16

OpenStudy (anonymous):

(x-1) = 8 x = 9.

jimthompson5910 (jim_thompson5910):

-2(x-1)=-16 -2x+2=-16 -2x=-16-2 -2x=-18 x=(-18)/(-2) x=9

OpenStudy (anonymous):

thanks.. now i see what i did wrong